# Distributional Diffusion Models with Scoring Rules

### v2

@@ Abstract
Diffusion models generate high-quality synthetic data. They operate by defining a continuous-time forward process which gradually adds Gaussian noise to data until fully corrupted. The corresponding reverse process progressively “denoises” a Gaussian sample into a sample from the data distribution. However, generating high-quality outputs requires many discretization steps to obtain a faithful approximation of the reverse process. This is expensive and has motivated the development of many acceleration methods. We propose to accomplish sample generation by learning the posterior distribution of clean data samples given their noisy versions, instead of only the mean of this distribution. This allows us to sample from the probability transitions of the reverse process on a coarse time scale, significantly accelerating inference with minimal degradation of the quality of the output. This is accomplished by replacing the standard regression loss used to estimate conditional means with a scoring rule. We validate our method on image and robot trajectory generation, where we consistently outperform standard diffusion models at few discretization steps.
@@ 1 Introduction
We depart from these earlier approaches in simply proposing to sample from the generative process on a coarser time scale. However, naively using denoisers obtained from the training of diffusion models significantly degrades the quality of the outputs, as these denoisers do not capture the full posterior distribution of the clean data given its noisy version, but rather its conditional mean. We review the relevant background on diffusions and their limitations in LABEL:{sec:diffusion_models}. In earlier works, it has been proposed by Xiao et al. (2022); Xu et al. (2024) to use Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) to learn such conditional distributions. We show here how to bypass adversarial training by relying on generalized scoring rules (Gneiting & Raftery, 2007), which we describe in Section 2.2. This yields a simple loss which interpolates between the classical regression loss used in diffusion models and distributional losses. We present our new class of Distributional Diffusion Models in Section 3, and provide theoretical grounding for our choice of distributional loss, as well as details of the interpolation to the regression loss, in Section 4. We review related work in Section 5.
In our experimental evaluation of Section 6, we demonstrate that our approach allows us to produce high-quality samples with significantly fewer denoising steps. We observe substantial benefits across a range of tasks, including image-generation tasks in pixel space and latent space, and in robotics applications. All proofs are given in the supplementary material and a table of notation is included in Appendix A for easy reference.
@@ 2.1 Diffusion models
It can be easily checked that this process has indeed the same marginal distributions, denoted $p_{t_{k}}(x_{t_{k}})$, as the process defined by $p_{0}(x_{t_{0}})p(x_{t_{1}:t_{N}}|x_{t_{0}})$ (see (1)) so that in particular $X_{t_{0}}\sim p_{0}$; see e.g. (Shi et al., 2024, Appendix E).
@@
for a weighting function $w_{t}$ (Kingma et al., 2021). Approximating $p_{0|t}(x_{0}|x_{t})$ with a Dirac mass located at $\hat{x}_{\theta}(t,x_{t})$ seems crude at first. However, as $N\rightarrow\infty$ with $t_{k+1}\to t_{k}$, the resulting discrete time process converges to a continuous time one which recovers the data distribution if $\hat{x}_{\theta}(t,X_{t})=\mathbb{E}[X_{0}|X_{t}]$; see e.g. (Song et al., 2021b).
When performing a coarse time discretization, this result no longer holds, and the Dirac approximation is poor. We propose here to learn a generative network to sample approximately from $p_{0|t}(x_{0}|x_{t})$.
@@ 2.2 Scoring rules
We next recall the framework of scoring rules, as reviewed by Gneiting & Raftery (2007), and then describe the specific scoring rules used in this work. Consider two probability distributions $p,q$. A scoring rule $S(p,x)$ indicates the quality of a prediction $p$ when the event $Y=y$ is observed. The expected score is its expectation under $q$,
@@
where $\rho$ is a continuous negative definite kernel. Similarly to Bouchacourt et al. (2016, Section 3.2), we will also consider generalized kernel scores of the form
@@
Positive definite kernels are said to be characteristic to $\mathcal{P}$ when $S_{\rho}$ yields a strictly proper scoring rule on $\mathcal{P}$ (Sriperumbudur et al., 2010, 2011). While the exponentiated quadratic (rbf) kernel satisfies this property, we will also investigate other kernels such as the inverse multiquadratic (imq) kernel and the exponential kernel (exp),
@@
When $S=S_{\rho}$ and $\rho(x,x^{\prime})=-k(x,x^{\prime})$ with $k$ a positive definite kernel, we recover the squared Maximum Mean Discrepancy ($\mathrm{MMD}^{2}$) (Gretton et al., 2012). For $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$, we obtain the energy distance (Rizzo & Székely, 2016; Sejdinovic et al., 2013).
@@ 3 Distributional Diffusion Models
Energy diffusion loss. Our final expected loss on $\theta$ integrates the conditional generalized energy score over both the noise level $t\in[0,1],$ and the samples from the dataset, $X_{0}\sim p_{0}$, together with noisy samples $X_{t}\sim p_{t|0}(\cdot|X_{0})$ so that $(X_{0},X_{t})\sim p_{0,t}$. This gives the energy diffusion loss
@@
where $w_{t}$ is a user-defined weighting function. The loss (14) has the following remarkable property: when we set $\beta\to 2$ and $\lambda\to 0$, we recover the classical regression diffusion loss (8). When $\beta\in(0,2)$ and $\lambda=1$ we obtain an (integrated) strictly proper expected conditional energy score. By selecting $\lambda\in(0,1)$ and $\beta\in(0,2]$, we can interpolate between these two cases. Our approach shares with (Bouchacourt et al., 2016; Gritsenko et al., 2020; Chen et al., 2024b; Pacchiardi et al., 2024; Shen & Meinshausen, 2024) the conditional energy score (13), but differs in that these earlier works learned only a single conditional distribution, whereas we integrate over many different noise levels.”
Empirical energy diffusion loss. For $n$ samples $\{X^{i}_{0}\}_{i=1}^{n}\overset{\mathrm{i.i.d.}}{\sim}p_{0}$ from the training dataset and corresponding $n$ noise levels $\{t_{i}\}_{i=1}^{n}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{U}[0,1]$, we sample noisy data points $X_{t_{i}}^{i}|X_{0}^{i},t_{i}\sim p_{t_{i}|0}(\cdot|X_{0}^{i})$ from the forward diffusion process (2) with noise schedule (3). Then, for each pair $(X_{0}^{i},X_{t_{i}}^{i})$, we sample $m$ Gaussian noise samples $\xi^{j}_{i}\sim\mathcal{N}(0,\mathrm{Id})$, $j\in[m]$ which allows us to compute $\hat{x}_{\theta}(t_{i},X_{t},\xi_{i}^{j})$. We then obtain the following empirical energy diffusion loss
$\mathcal{L}_{n,m}(\theta){=\frac{1}{nm}\sum_{i,j=1}^{n,m}w_{t_{i}}\Bigg{[}\left\|X_{0}^{i}-\hat{x}_{\theta}(t_{i},X_{t},\xi_{i}^{j})\right\|^{\beta}}$ (15)
@@
Figure 1: Left: reduction factor $f(\lambda,\beta)$ as a function of $\beta\in[0,2]$ and $\lambda\in[0,1]$. Right: evolution of the covariance during the sampling of distributional diffusion model for a Gaussian target $p_{0}=\mathcal{N}(0,4\mathrm{Id})$ for $20$ of $N\in\{5,10,\dots,95,100\}$, $\lambda=0.5$. As $N$ increases, the output variance gets closer to $4$.
@@
Sampling. Once we have trained our model $\hat{x}_{\theta}$, we generate samples using Algorithm 2. This procedure is very similar to DDIM (Song et al., 2021a). The only difference is that now $\hat{x}_{\theta}$ outputs an approximate sample from $p_{0|t}(x_{0}|x_{t})$ instead of an approximation of $\mathbb{E}[X_{0}|X_{t}=x_{t}]$.
@@ 4.1 Gaussian analysis
We shed more light on the two main hyperparameters of the generalized kernel score introduced in Section 2.2 (i) the kernel parameter ($\beta$ in the case of the energy score) (ii) the trade-off term $\lambda$. We focus here on the case of generalized energy score defined in (2.2).
@@
A notable outcome of maximizing the generalized energy score is that $\mu^{\star}=\mu$ no matter the value of $\lambda$ and $\beta$. With $\lambda=1$ and $\beta<2$, we also recover the correct variance, which is to be expected as $S_{\lambda,\beta}$ is a strictly proper energy score. As soon as $\lambda<1$, the variance $\sigma^{2}$ is underestimated by a factor $f(\lambda,\beta)=(2\lambda^{-2/(2-\beta)}-1)^{-1}$; see Figure 1.
We now consider the effect of these hyperparameters when sampling from the corresponding generative model using Algorithm 2. For simplicity, we let the churn parameter $\varepsilon=1$ but our analysis also applies to $\varepsilon\in[0,1)$. We consider a Gaussian target $p_{0}=\mathcal{N}(\mu,\sigma^{2}\mathrm{Id})$. In that case, $p_{0|t}(x_{0}|x_{t})$ is also Gaussian for any $t\in[0,1]$. Therefore if this density was estimated using the generalized energy score, the results of Section 4.1 apply. In particular, this implies that the induced estimate of the Gaussian distribution $p_{s|t}$ obtained by plugging our approximation of $p_{0|t}$ into (7) would have the correct mean but an incorrect variance for $\lambda<1$.
As expected, differences between classical diffusion models and distributional diffusion models arise when considering larger discretization stepsizes, where the Dirac approximation of $p_{0|t}(x_{0}|x_{t})$ is no longer valid, see Figure 1.
@@ 5 Related Work
Discrepancy and diffusion. We modify the diffusion model training loss to learn the conditional distribution $p_{0|t}(x_{0}|x_{t})$. The importance of approximating the covariance this distribution was noted in (Nichol & Dhariwal, 2021) and exploited in (Ho et al., 2020; Nichol & Dhariwal, 2021; Bao et al., 2022a, b; Ou et al., 2024). Closest to our approach is (Xiao et al., 2022), which uses a GAN to better learn $p_{0|t}(x_{0}|x_{t})$ for fewer sampling steps. While sharing the same motivation, our method avoids adversarial training and discriminator training, offering a simple loss modification that encompasses standard diffusion models. Other loss modifications, like using the $\ell_{1}$ loss (Chen et al., 2021; Saharia et al., 2022a), aim to improve output quality, not reduce sampling steps. (Galashov et al., 2025) can be seen as dual to our work: they learn discriminative kernel features with a flow at different noise levels, while we focus solely on learning a generator.
@@ 6 Experiments
In this section, we validate the performance of our approach in 2D, image generation, and robotics settings. The main experiments are presented here, while additional experiments and ablations are given in Appendix J. All the experimental details are described in Appendix I.
@@ 6.2 Image experiments
Main results. We train conditional pixel-space models on CIFAR-10 (32x32x3) and on CelebA (64x64x3), as well as unconditional pixel-space models on LSUN Bedrooms (64x64x3). We further use an autoencoder trained on CelebA-HQ (256x256x3) producing latent codes of shape 64x64x3, which are then used to build latent unconditional models. For each dataset, we train a diffusion model by optimizing (8) as well as a distributional model (Algorithm 1) for different $\lambda\in[0,1]$ and $\beta\in[0.001,2]$. Models are evaluated using the FID (Heusel et al., 2017) score. See Section I.2 for details.
@@
In Figure 5, we show results of conditional image generation on CIFAR-10 and CelebA, and in Figure 6 results of unconditional image generation on LSUN Bedrooms and latent CelebA-HQ. Whenever we report performance of distributional model and write $\beta<2$, we select a parameter $\beta$ given one fixed $\lambda$ which minimizes the FID. More detailed results are given in Section J.1, which include FIDs for different numbers of steps for every combination of $(\beta,\lambda)$.
Main takeaways. The results suggest that whenever the number of diffusion steps is low, then distributional models with $\lambda=1,\beta<2$ achieve better performance than classical diffusions. This confirms that the diffusion model approximation is poor in the ”few steps” regime, and can be improved with better modeling of $p_{0|t}$ as explained in Section 2.1. As the number of diffusion steps increases, the Dirac approximation becomes more accurate and diffusion models achieve the best performance. In that scenario, we observe that distributional models with $\lambda=1,\beta<2$ yield worse performance. Our hypothesis is that since distributional models $\hat{x}_{\theta}$ only approximate the posterior distribution $p_{0|t}$, this could lead to an increased amount of noise at sampling time and consequently to error accumulation. However, we observe that distributional model variants with $\lambda\in(0,1)$ still yield good performance as the number of diffusion steps increases. This highlights the trade-off between distributional and diffusion models controlled by the parameter $\lambda$. As argued in Section 4.1, $\lambda<1$ leads to underestimation of the target variance which, for denoising diffusions, implies that the variance of $\hat{x}_{\theta}$ is more concentrated around $p_{0|t}$ than would occur for posterior samples. Remarkably, we also found that using $\beta<2,\lambda=0$ led to comparable or even slightly better performance compared to classical diffusion models. This suggests that it could be valuable to train diffusion models using a different loss than Mean Square Error (MSE) regression.
@@
Conditional vs unconditional generation. We observe slightly different behavior of distributional models in the conditional image generation case (Figure 5), and in the unconditional case (Figure 6). For a small number of diffusion steps, we notice that the gap in performance between the model trained with $\lambda=1,\beta<2$ and a diffusion model, is much smaller in the unconditional case (1.5x) compared to the conditional one (3x). In Figure 7 we visualize the best parameters $\lambda,\beta$ selected to achieve the lowest FID, as a function of diffusion steps. More detailed results are presented in Section J.1. We observe that overall $\lambda$ follows a downward trend, decreasing as the number of steps increases. However, $\beta$ behaves differently in each setting, increasing monotonically as a function of diffusion steps in the conditional case, while remaining at a minimal value in the unconditional case. We hypothesise that this is because learning unconditional distributions is a much harder task since the data is spread out more, leading to a much higher variance in the distribution of $\hat{x}_{\theta}$.
Using different kernels. As discussed in Section 2.2 and in Section 3, we can also use the kernel diffusion loss. We train $\hat{x}_{\theta}$ with kernels (9a), (9b), (9c), varying kernel parameters and $\lambda\in[0,1]$; see Figure 8 for results on conditional image generation for CIFAR-10. We observe similar behavior to the energy diffusion loss, achieving the best results when $\lambda=1$ for few diffusion steps and better performance with $\lambda<1$ when number of steps increases; see Section J.1 for detailed results.

### v3

@@ Abstract
Diffusion models generate high-quality synthetic data. They operate by defining a continuous-time forward process which gradually adds Gaussian noise to data until fully corrupted. The corresponding reverse process progressively “denoises” a Gaussian sample into a sample from the data distribution. However, generating high-quality outputs requires many discretization steps to obtain a faithful approximation of the reverse process. This is expensive and has motivated the development of many acceleration methods. We propose to speed up sample generation by learning the posterior distribution of clean data samples given their noisy versions, instead of only the mean of this distribution. This allows us to sample from the probability transitions of the reverse process on a coarse time scale, significantly accelerating inference with minimal degradation of the quality of the output. This is accomplished by replacing the standard regression loss used to estimate conditional means with a scoring rule. We validate our method on image and robot trajectory generation, where we consistently outperform standard diffusion models at few discretization steps.
@@ 1 Introduction
We depart from these earlier approaches in simply proposing to sample from the generative process on a coarser time scale. However, naively using denoisers obtained from the training of diffusion models significantly degrades the quality of the outputs, as these denoisers do not capture the full posterior distribution of the clean data given its noisy version, but rather its conditional mean. We review the relevant background on diffusions and their limitations in LABEL:{sec:diffusion_models}. In earlier works, Xiao et al. (2022); Xu et al. (2024) proposed to use Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) to learn such conditional distributions. We show here how to bypass adversarial training by relying on generalized scoring rules (Gneiting & Raftery, 2007), which we describe in Section 2.2. This yields a simple loss which interpolates between the classical regression loss used in diffusion models and distributional losses. We present our new class of Distributional Diffusion Models in Section 3, and provide theoretical grounding for our choice of distributional loss, as well as details of the interpolation to the regression loss, in Section 4. We review related work in Section 5.
Experiments in Section 6 demonstrate that our approach produces high-quality samples with significantly fewer denoising steps. We observe substantial benefits across a range of tasks, including image-generation tasks in both pixel and latent spaces, and in robotics applications. All proofs are in the supplementary and a table of notation is in Appendix A.
@@ 2.1 Diffusion models
It can be easily checked that this process has the same marginal distributions, denoted $p_{t_{k}}(x_{t_{k}})$, as the process defined by $p_{0}(x_{t_{0}})p(x_{t_{1}:t_{N}}|x_{t_{0}})$ (see (1)) so that in particular $X_{t_{0}}\sim p_{0}$; see e.g. (Shi et al., 2024, Appendix E).
@@
for a weighting function $w_{t}$ (Kingma et al., 2021). Approximating $p_{0|t}(x_{0}|x_{t})$ with a Dirac mass located at $\hat{x}_{\theta}(t,x_{t})$ seems crude at first. However, as $N\rightarrow\infty$ with $t_{k+1}\to t_{k}$, the resulting discrete time process converges to a continuous time process which recovers the data distribution if $\hat{x}_{\theta}(t,X_{t})=\mathbb{E}[X_{0}|X_{t}]$; see e.g. Song et al. (2021b).
When performing a coarse time discretization, however, this result no longer holds, and the Dirac approximation is poor. We propose here to learn a generative network to sample approximately from $p_{0|t}(x_{0}|x_{t})$.
@@ 2.2 Scoring rules
We next recall the framework of scoring rules, as reviewed by Gneiting & Raftery (2007), and then describe the specific scoring rules used in this work. Consider two probability distributions $p,q$. A scoring rule $S(p,y)$ indicates the quality of a prediction $p$ when the event $Y=y$ is observed. The expected score is its expectation under $q$,
@@
where $\rho$ is a continuous negative definite kernel. Following Bouchacourt et al. (2016, Section 3.2), we will also consider generalized kernel scores of the form
@@
Positive definite kernels are said to be characteristic to $\mathcal{P}$ when $S_{\rho}$ yields a strictly proper scoring rule on $\mathcal{P}$ (Sriperumbudur et al., 2010, 2011). While the exponentiated quadratic (rbf) kernel satisfies this property, we will also investigate other kernels such as the inverse multiquadratic (imq) kernel and the exponential (exp) kernel,
@@
When $S=S_{\rho}$ and $\rho(x,x^{\prime})=-k(x,x^{\prime})$ with $k$ a positive definite kernel, we recover the squared Maximum Mean Discrepancy ($\mathrm{MMD}^{2}$) (Gretton et al., 2012). For $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$, we obtain the energy distance (Rizzo & Székely, 2016). The relation between these discrepancies was established by Sejdinovic et al. (2013) where the specific positive definite kernel family for which the MMD and energy distance are equivalent is described.
@@ 3 Distributional Diffusion Models
Energy diffusion loss. Our final expected loss integrates the conditional generalized energy score over both the noise level $t\in[0,1],$ and the samples from the dataset, $X_{0}\sim p_{0}$, together with noisy samples $X_{t}\sim p_{t|0}(\cdot|X_{0})$ so that $(X_{0},X_{t})\sim p_{0,t}$. This gives the energy diffusion loss
@@
where $w_{t}$ is a user-defined weighting function. The loss (14) has the following remarkable property: when we set $\beta\to 2$ and $\lambda\to 0$, we recover the classical regression diffusion loss (8). When $\beta\in(0,2)$ and $\lambda=1$ we obtain an (integrated) strictly proper expected conditional energy score. By selecting $\lambda\in(0,1)$ and $\beta\in(0,2]$, we can interpolate between these two cases. Our approach shares with (Bouchacourt et al., 2016; Gritsenko et al., 2020; Chen et al., 2024b; Pacchiardi et al., 2024; Shen & Meinshausen, 2024) the conditional energy score (13), but differs in that these earlier works learned only a single conditional distribution, whereas we integrate over many different noise levels.
Empirical energy diffusion loss. For samples $\{X^{i}_{0}\}_{i=1}^{n}\overset{\mathrm{i.i.d.}}{\sim}p_{0}$ from the training set and corresponding noise levels $\{t_{i}\}_{i=1}^{n}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{U}[0,1]$, we sample noisy data points $X_{t_{i}}^{i}|X_{0}^{i},t_{i}\sim p_{t_{i}|0}(\cdot|X_{0}^{i})$ from the forward diffusion process (2) with noise schedule (3). Then, for each pair $(X_{0}^{i},X_{t_{i}}^{i})$, we sample $m$ Gaussian noise samples $\xi^{j}_{i}\sim\mathcal{N}(0,\mathrm{Id})$, $j\in[m]$ which allows us to compute $\hat{x}_{\theta}(t_{i},X_{t_{i}},\xi_{i}^{j})$. We then obtain the following empirical energy diffusion loss
$\mathcal{L}_{n,m}(\theta){=\frac{1}{nm}\sum_{i,j=1}^{n,m}w_{t_{i}}\Bigg{[}\left\|X_{0}^{i}-\hat{x}_{\theta}(t_{i},X_{t_{i}},\xi_{i}^{j})\right\|^{\beta}}$ (15)
@@
Figure 1: Left: variance reduction factor $f(\lambda,\beta)$ as a function of $\beta\in[0,2]$ and $\lambda\in[0,1]$. Right: evolution of the covariance during the sampling of distributional diffusion model for a Gaussian target $p_{0}=\mathcal{N}(0,4\mathrm{Id})$ with a number of denoising steps $N\in\{5,10,\dots,95,100\}$, $\lambda=0.5,\beta=0.2$. As $N$ increases (the color changes from blue to red), the output variance gets closer to $4$.
@@
Sampling. Once we have trained our model $\hat{x}_{\theta}$, we generate samples using Algorithm 2. This procedure is very similar to DDIM (Song et al., 2021a). The only difference is that $\hat{x}_{\theta}$ now outputs an approximate sample from $p_{0|t}(x_{0}|x_{t})$ instead of an approximation of $\mathbb{E}[X_{0}|X_{t}=x_{t}]$.
@@ 4.1 Gaussian analysis
We shed more light on the two main hyperparameters of the generalized kernel score introduced in Section 2.2: (i) the kernel parameter ($\beta$ in the case of the energy score); (ii) the trade-off term $\lambda$. We focus here on the case of generalized energy score defined in (2.2).
@@
A notable outcome of maximizing the generalized energy score is that $\mu^{\star}=\mu$ no matter the value of $\lambda$ and $\beta$. With $\lambda=1$ and $\beta<2$, we recover as expected the correct variance, since $S_{\lambda,\beta}$ is a strictly proper energy score. As soon as $\lambda<1$, the variance $\sigma^{2}$ is underestimated by a factor $f(\lambda,\beta)=(2\lambda^{-2/(2-\beta)}-1)^{-1}$; see Figure 1, left.
We now consider the effect of these hyperparameters when sampling from the corresponding generative model using Algorithm 2. For simplicity, we let the churn parameter $\varepsilon=1$ but our analysis also applies to $\varepsilon\in[0,1)$. We consider a Gaussian target $p_{0}=\mathcal{N}(\mu,\sigma^{2}\mathrm{Id})$. In that case, $p_{0|t}(x_{0}|x_{t})$ is also Gaussian for any $t\in[0,1]$. Therefore if this density were estimated using the generalized energy score, then the results of Section 4.1 apply. In particular, this implies that the induced estimate of the Gaussian distribution $p_{s|t}$ obtained by plugging our approximation of $p_{0|t}$ into (7) would have the correct mean but an incorrect variance for $\lambda<1$.
As expected, differences between classical diffusion models and distributional diffusion models arise when considering larger discretization stepsizes, where the Dirac approximation of $p_{0|t}(x_{0}|x_{t})$ is no longer valid, see Figure 1, right.
@@ 5 Related Work
Discrepancy and diffusion. We modify the diffusion model training loss to learn the conditional distribution $p_{0|t}(x_{0}|x_{t})$. The importance of approximating the covariance of this distribution was noted by Nichol & Dhariwal (2021) and exploited in (Ho et al., 2020; Nichol & Dhariwal, 2021; Bao et al., 2022a, b; Ou et al., 2024). Closest to our approach is (Xiao et al., 2022), which uses a GAN to learn $p_{0|t}(x_{0}|x_{t})$ for fewer sampling steps. While sharing the same motivation, our method avoids adversarial training and discriminator training, offering a simple loss modification that encompasses standard diffusion models. Other loss modifications, like using the $\ell_{1}$ loss (Chen et al., 2021; Saharia et al., 2022a), aim to improve output quality, not reduce sampling steps. Galashov et al. (2025) can be seen as dual to our work: they learn discriminative kernel features with a flow at different noise levels, while we focus solely on learning a generator.
@@ 6 Experiments
In this section, we validate the performance of our approach in 2D, image generation, and robotics settings. The main experiments are presented here, while additional experiments and ablations are given in Appendix K. All the experimental details are described in Appendix I, while Appendix J analyzes computational complexity.
@@ 6.2 Image experiments
Main results. We train conditional pixel-space models on CIFAR-10 (32x32x3) and on CelebA (64x64x3), as well as unconditional pixel-space models on LSUN Bedrooms (64x64x3). We further use an autoencoder trained on CelebA-HQ (256x256x3) producing latent codes of shape 64x64x3, which are then used to build latent unconditional models. For each dataset, we train a diffusion model by optimizing (8) as well as a distributional model (Algorithm 1) for different $\lambda\in[0,1]$ and $\beta\in[0.001,2]$. Models are evaluated using the FID score (Heusel et al., 2017). See Section I.2 for details.
@@
In Figure 5, we show results of conditional image generation on CIFAR-10 and CelebA, and in Figure 6 results of unconditional image generation on LSUN Bedrooms and latent CelebA-HQ. Whenever we report performance of distributional model and write $\beta<2$, we select a parameter $\beta$ given one fixed $\lambda$ which minimizes the FID. More detailed results are given in Section K.2, which include FIDs for different numbers of steps for every combination of $(\beta,\lambda)$.
Main takeaways. The results suggest that whenever the number of diffusion steps is low, then distributional models with $\lambda=1,\beta<2$ achieve better performance than classical diffusion models. This confirms that the diffusion model approximation is poor in the ”few steps” regime, and can be improved with better modeling of $p_{0|t}$ as explained in Section 2.1. As the number of diffusion steps increases, the Dirac approximation becomes more accurate and diffusion models achieve the best performance. In that scenario, we observe that distributional models with $\lambda=1,\beta<2$ yield worse performance. Our hypothesis is that since distributional models $\hat{x}_{\theta}$ only approximate the posterior distribution $p_{0|t}$, this could lead to an increased amount of noise at sampling time and consequently to error accumulation. However, we observe that distributional model variants with $\lambda\in(0,1)$ still yield good performance as the number of diffusion steps increases. This highlights the trade-off between distributional and diffusion models controlled by the parameter $\lambda$. As argued in Section 4.1, $\lambda<1$ leads to underestimation of the target variance which, for denoising diffusions, implies that the variance of $\hat{x}_{\theta}$ is more concentrated around $p_{0|t}$ than would occur for posterior samples. Remarkably, we also found that using $\beta<2,\lambda=0$ led to comparable or even slightly better performance compared to classical diffusion models. This suggests that it could be valuable to train diffusion models using a different loss than Mean Square Error (MSE) regression.
@@
Conditional vs unconditional generation. We observe slightly different behavior of distributional models in the conditional image generation case (Figure 5), and in the unconditional case (Figure 6). For a small number of diffusion steps, we notice that the gap in performance between the model trained with $\lambda=1,\beta<2$ and a diffusion model, is much smaller in the unconditional case (1.5x) compared to the conditional one (3x). In Figure 7 we visualize the best parameters $\lambda,\beta$ selected to achieve the lowest FID, as a function of diffusion steps. More detailed results are presented in Section K.2. We observe that overall $\lambda$ follows a downward trend, decreasing as the number of steps increases. However, $\beta$ behaves differently in each setting, increasing monotonically as a function of diffusion steps in the conditional case, while remaining at a minimal value in the unconditional case. We hypothesise that this is because learning unconditional distributions is a much harder task since the data is spread out more, leading to a much higher variance in the distribution of $\hat{x}_{\theta}$.
Using different kernels. As discussed in Section 2.2 and in Section 3, we can also use the kernel diffusion loss. We train $\hat{x}_{\theta}$ with kernels (9a), (9b), (9c), varying kernel parameters and $\lambda\in[0,1]$; see Figure 8 for results on conditional image generation for CIFAR-10. We observe similar behavior to the energy diffusion loss, achieving the best results when $\lambda=1$ for few diffusion steps and better performance with $\lambda<1$ when number of steps increases; see Section K.2 for detailed results.
@@
Comparisons to distillation and numerical solvers. In Section K.3, we compare our approach with the DPM-solver++ method Lu et al. (2022) and the state-of-the-art, multi-step distillation method, moment-matching distillation (Salimans et al., 2024). Overall, our approach achieves results competitive with multistep distillation (where we compared different numbers of steps) and outperforms DPM-solver++ for more than 8 NFEs. The performance of DPM-solver++ degrades as the number of steps increases, which is expected due to its numerical instability.
Our approach offers two compelling benefits over multistep distillation. First, it does not rely on distillation, eliminating the need to train a large (and slow) teacher model. Second, it does not require specifying additional sampler hyperparameters during training. Combining our distributional approach with a modern distillation method is an interesting direction for future research.